The sum of two numbers is 38. Their difference is 8. What are the two numbers?
Answer:
First number:
Second number

Answer by stanbon(75887)

The sum of two numbers is 38. Their difference is 8. What are the two numbers?
--------
EQUATIONS:
x+y=38
x-y=8
----------
Add the two fractions to get:
2x=46
x=23
--------
Substitute to solve for y:
23-y=8
y=23-8
y=15
============
Cheers,
Stan H.